Sharon is three times as old as her daughter. If Sharon is 32 years older than her daughter, what are their ages?

Answer by checkley79(3341)

S=3D
S=D+32
3D=D+32
3D-D=32
2D=32
D=32/2
D=16 YEARS OLD FOR THE DAUGHTER.
PROOF:
3*16=16+32
48=48